OpenStudy (anonymous):

if x=cosh theta, y=sinh theta, obtain dy/dx in parameter theta. then sketch graph of y and dy/dx regarded as function theta

OpenStudy (irishboy123):

\[\dfrac{d (\sinh \theta)}{d (\cosh \theta)}= \dfrac{y'_\theta}{x'_\theta} = \coth \theta\] and then play with \( csch ^{2} x= \coth^{2} x - 1\)

OpenStudy (astrophysics):

Nice

OpenStudy (anonymous):

the identity is use to obtain what actually?

OpenStudy (irishboy123):

you wanna flip it back into x and y, i think

OpenStudy (anonymous):

i see, thank you for your help :)

OpenStudy (irishboy123):

mp